OpenStudy (anonymous):

how do i find inverse of sqrt x +4 divided by 7-sqrt of x ??

OpenStudy (anonymous):

Is this the expression you want to find the inverse of: \(\Large y=\frac{\sqrt{x+4}}{7-\sqrt{x}} \) ????

OpenStudy (anonymous):

If yes, then you'll need to solve for y here: \(\Large x=\frac{\sqrt{y+4}}{7-\sqrt{y}} \) The expression you'll get for y represents the inverse.